# Date of publication: December 8th, 2025

### v1

@@ VII-C1 Hardware integration of the ground system
  • Install 7 smart DC meters (solar + DC load). Meter accuracy must be Class-0.5 or higher.
  • Install 3 smart AC meters (220 VAC output). Meter accuracy must be Class-0.5 or higher.
  • Install 1 smart AC meter (110 VAC output). Meter accuracy must be Class-0.5 or higher.
  • Add 4 industrial-grade embedded systems (for input/output). Hardware-wise, it must have an industrial-grade storage temperature range (-40 80C) and an industrial-grade operating temperature range (-20 60C) to cope with high-altitude climate changes, and possess communication interfaces such as Ethernet, UART, RS-485, RS-232, and CAN Bus. Software-wise, it adopts a Linux-based operating system architecture, capable of decoding and performing input/output operations for Ethernet, UART, RS-485, RS-232, and CAN Bus communications. For storage, it uses industrial-grade, high-durability SD cards.
@@ VII-C3 Software integration of the cloud system
  • Computing units must be hosted on a reliable and scalable infrastructure. A Service-Level Agreement (SLA) commitment with 99.99% availability is required, with capacity scalable within minutes. A secure computing environment must also be provided.
@@
  • The data storage service used must be scalable, data available, secure, and efficient. It should protect any amount of data, such as data lakes, websites, cloud-native applications, backups, archives, machine learning, and analytics. Its storage service should be designed to achieve 99.999999999% (11 nines) durability.
@@ VII-D AC and DC systems integration
This project involves both AC and DC systems within a microgrid. The power lines for these two systems must be independent; otherwise, system damage may occur. Appropriate physical isolation (e.g., breakers, switches, etc.) is required between each device to facilitate installation and subsequent maintenance. Other points to note:
@@
    The inverter’s DC operating voltage setting must match the operating voltage setting of the LFP ESS to prevent the inverter from draining the battery’s power.
@@
Junction boxes are divided into three main categories: PV panel, AC panel, and DC panel. The PV panel provides DC power transmission between the PV arrays and hybrid solar inverters. The AC panel provides power transmission from the hybrid solar inverters to 220 VAC loads, from the generator to the hybrid solar inverters, and from the 220 VAC power to the AC transformer, converting the 220 VAC power to 110 VAC to supply 110 VAC loads. The DC panel provides bidirectional power transmission between the hybrid solar inverters and the LFP ESS.
Additionally, a self-developed cloud-native EMS can detect PV power generation (W), PV energy generation (kWh), load AC power consumption (W), and load AC energy consumption (kWh), showing the values on the enclosure using meters with LCD. The information is also uploaded to a cloud server for remote monitoring.
@@ VIII Summary and Outlook
In the next step, the remaining lead-acid batteries should be replaced by the eco-friendly LFP ESS entirely, making a better and greener off-grid solar ESS. On the other hand, a more comprehensive EMS should be installed to check the energy input and output. Figuring out the energy usage of the cabins. Further suggestions can be given after the energy usage analysis, such as “Is it necessary to expand the PV system?” or “Is it necessary to expand the ESS?”

### v2

@@ VII-C1 Hardware integration of the ground system
  • Install 7 smart DC meters (solar + DC load). Meter accuracy should be Class-0.5 or higher.
  • Install 3 smart AC meters (220 VAC output). Meter accuracy should be Class-0.5 or higher.
  • Install 1 smart AC meter (110 VAC output). Meter accuracy should be Class-0.5 or higher.
[…]
@@ VII-C3 Software integration of the cloud system
  • Computing units should be hosted on a reliable and scalable infrastructure. A Service-Level Agreement (SLA) commitment with 99.99% availability is required, with capacity scalable within minutes. A secure computing environment should also be provided.
@@
  • The data storage service used should be scalable, data available, secure, and efficient. It should protect any amount of data, such as data lakes, websites, cloud-native applications, backups, archives, machine learning, and analytics. Its storage service should be designed to achieve 99.999999999% (11 nines) durability.
@@ VII-D AC and DC systems integration
This project involves both AC and DC systems within a microgrid. The power lines for these two systems should be independent; otherwise, system damage may occur. Appropriate physical isolation (e.g., breakers, switches, etc.) is required between each device to facilitate installation and subsequent maintenance. Other points to note:
@@
    The inverter’s DC operating voltage setting should match the operating voltage setting of the LFP ESS to prevent the inverter from draining the battery’s power.
@@
Junction boxes are divided into three main categories: PV panel, AC panel, and DC panel. The PV panel provides DC power transmission between the PV arrays and hybrid solar inverters. The AC panel provides power transmission from the hybrid solar inverters to 220 VAC loads, from the generator to the hybrid solar inverters, and from the 220 VAC power to the AC transformer, converting the 220 VAC power to supply 110 VAC loads. The DC panel provides bidirectional power transmission between the hybrid solar inverters and the LFP ESS.
Additionally, a self-developed cloud-native EMS can detect PV power generation (unit: W), PV energy generation (unit: kWh), load AC power consumption (unit: W), and load AC energy consumption (unit: kWh), showing the values on the enclosure using meters with liquid crystal display (LCD). The information is also uploaded to a cloud server for remote monitoring.
@@ VIII Summary and Outlook
In the next step, the remaining lead-acid batteries should be replaced by the eco-friendly LFP ESS entirely, making a better and greener off-grid solar ESS. On the other hand, a more comprehensive EMS should be installed to check the energy input and output. Figuring out the energy usage of the cabins. Further suggestions can be given after the energy usage analysis, such as “Is the energy generated sufficient? Is it necessary to expand the PV system?” or “Is the energy stored sufficient? Is it necessary to expand the ESS?”